If the line x+y-1=0 touches parabola y^2=kx, then  k? 

Grade:

1 Answers

kkbisht
5 years ago
Hellow friends : The answer is here:
The line x+y=1 touches the parabola y^2=kx or y^2=4.k/4 x ( Compare with standard eqn y^2=4ax) so a= k/4 and the condition of tangency for the line y=mx+c  to a parabola y^2=4ax is that : c=a/m
we ghave given c=1,  m=-1, a=k/4 This implies 1=k/4(-1) => k=-4
 
k.k.Bisht